OpenStudy (anonymous):

Suppose the ice cream on top of a cone makes a perfectly spherical shape. If the diameter of the ice cream scoop was 6 inches, what will happen to the volume of the ice cream when the ice cream scoop is decreased to a diameter of 3 inches? What will the original and new volume of the ice cream be?

OpenStudy (austinl):

The formula for the volume of a sphere is, \[V=\frac{4}{3}(\pi)r^3\] Can you go from here?

OpenStudy (anonymous):

no

OpenStudy (austinl):

Ok, so the original diameter is 6 inches right? So what is the radius?

OpenStudy (anonymous):

V=4/3(3.14)3^3

OpenStudy (anonymous):

radius=3

OpenStudy (austinl):

Ok, good. So, you would plug that into the formula. Also, what class is this for?

OpenStudy (anonymous):

Integrated Math II Unit 8: Geometry of Spheres Section B: Investigating Changes in Measurement

OpenStudy (anonymous):

I got 113.04

OpenStudy (austinl):

Ok, so you are probably going to want to keep your answer in its exact form, or with pi still in the answer. For this you would take (4/3)(3^3), and then leave pi on the end of the answer. What do you get?

OpenStudy (anonymous):

Same answer 113.04

OpenStudy (austinl):

What is (4/3)(3^3)?

OpenStudy (anonymous):

volume

OpenStudy (anonymous):

It says what is the original and new volume though?

OpenStudy (austinl):

No, I mean literally what is (4/3) multiplied by (3^3)?

OpenStudy (anonymous):

36

OpenStudy (austinl):

Ok, so to leave it in its exact form, you would have \[36\pi\] That is the original Volume.

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

What about the new volume?

OpenStudy (austinl):

Ok, for the new volume. The diameter is 3, so what is the new radius?

OpenStudy (anonymous):

1.5

OpenStudy (austinl):

Ok, so plug that into the equation. What do you get?

OpenStudy (anonymous):

4/3(3.14)(1.5^3)

OpenStudy (anonymous):

14.3

OpenStudy (austinl):

Ok, be sure to leave it in terms of pi, it makes the last part easier. Take the original minus the new volume. In terms of pi, what do you get?

OpenStudy (anonymous):

98.74

OpenStudy (anonymous):

And what do you mean by terms of pi?

OpenStudy (austinl):

Ok, the first answer is \[36\pi\] The second part is \[(\frac{9}{2})\pi\]

OpenStudy (austinl):

That is what I mean by in terms of pi.

OpenStudy (anonymous):

Thx for trying

OpenStudy (anonymous):

But your answer is kinda wrong

OpenStudy (anonymous):

Got to go

OpenStudy (austinl):

Ok. Good luck, if you need help feel free to ask.